OpenStudy (anonymous):

Multiplication Property of Equality 5x/9==36 then, ____

OpenStudy (anonymous):

@PeachRings

OpenStudy (anonymous):

i dunno what else to say than 5x=36*9 x=36*9/5

OpenStudy (anonymous):

idk either but ima go with then 36 = 5x/9 that's one of my choices hat makes the most sense to me

OpenStudy (anonymous):

sounds like a plan